The reaction of 6.70 g of carbon with excess O2 yields 10.3 g of CO2. What is the percent yield of this reaction?
1
Expert's answer
2020-03-31T12:31:42-0400

C+O2=CO2

Ƞ=m/mt


m(CO2)=10.3g

n(C)=m(C)/M(C)=6.70/12=0.558mol

According to the equation of a chemical reaction n(CO2)=n(C)=0.558mol

mt(CO2)=n(CO2)*M(CO2)=0.558*(12+16*2)=0.558*44=24.57g

 

Ƞ=(m/mt)*100%=(10.3/24.57)*100%=41,9%

 

Answer: 41,9%